OpenStudy (anonymous):

radical 40 plus radical 100 how to do??

OpenStudy (amistre64):

you have to see if you can rewrite them with like radical parts

OpenStudy (anonymous):

...

OpenStudy (anonymous):

step by step?

OpenStudy (amistre64):

rewrite the given terms in a new way such that they each have the same radicals

OpenStudy (amistre64):

\[\sqrt{40}=\sqrt{4*10}=\sqrt{4}\sqrt{10}=2\sqrt{10}\]

OpenStudy (amistre64):

you can do the same methods for the 100

OpenStudy (amistre64):

maybe

OpenStudy (anonymous):

im still confused...

OpenStudy (amistre64):

really? sqrt(100) = 10 :)

OpenStudy (amistre64):

theres no good way to combine them i spose

OpenStudy (amistre64):

2sqrt(10) + 10 is the best you can get if im seeing it right

OpenStudy (anonymous):

im gonna ask someone else, thanks anyway.